Okay: That question is- The width of a dock is 4m less than the length. The area is 12m squared.
So i know that I am trying to find the length so that obviously is my x. Then i also know that the width is 4m less than that, but i get so thrown off by the 12 squared that i can't figure out how to set up the problem :/

Answer by richwmiller(17219)

What is the area ? length times width.
You actually don't say what we are looking for. Are we looking for the width, length or both?
I will use A for area L for length and W for width
L*W=A
L*W=12
W=L-4
L*(L-4)=12
L^2-4L-12=0
(L-6)(L+2)=0
L=6 meters
W=L-4=2 meters